# Parallel Diffusion Solver via Residual Dirichlet Policy Optimization

### v1

@@ III-B1 Discussion with multi-step solvers
While multi-step solvers [lu2022dpm_plus, zhao2024unipc, liupseudo, zhangfast] also combine multiple gradients to approximate the integral, they fundamentally differ in where these gradients are evaluated. Specifically, multi-step methods rely on Taylor expansion or polynomial extrapolation to linearly combine historical gradients evaluated at previous time steps, i.e., outside the current integration interval. In contrast, Theorem 1 implies that the exact integral over an interval $[a,b]$ admits a representation as a convex combination of gradients evaluated at points strictly within $(a,b)$. Motivated by this theoretical result, our method explicitly constructs a simplex-weighted combination of multiple gradients evaluated within the current time interval, leading to a more faithful approximation of the integral.
@@ III-B2 Discussion with AMED-Solver
AMED-Solver [zhou2024fast] estimates the update direction using a single intermediate timestep. However, according to Theorem 1, the integral of a vector-valued function cannot, in general, be exactly represented by the derivative at a single timestep; instead, it admits a convex combination of derivatives evaluated at multiple timesteps. A single timestep may only suffice when the underlying trajectory is effectively one-dimensional. In the sequel, we empirically analyze the geometric properties of diffusion sampling trajectories and find that they are nearly confined to a two-dimensional manifold, violating this condition.
In Figure 3, we analyze diffusion sampling trajectories using the models from [Karras2024edm2]. Despite the high dimensionality of the ambient space, the cumulative explained variance ratio shows that over 97% of the residual variance is captured by the first two principal components, indicating that the non-linear trajectory is effectively confined to a two-dimensional manifold. As a result, single-step or single-intermediate-point methods such as AMED-Solver are generally insufficient to characterize the local curvature within this plane using a single direction. This motivates the use of multiple intermediate gradients to span the underlying subspace and achieve a more accurate integral approximation.
@@ III-C Stage 1: Distillation-Based Parameter Optimization
[ningelucidating, li2024alleviating] identify exposure bias—i.e., the mismatch between training and sampling inputs—as a key factor contributing to error accumulation and sampling drift. To mitigate this, they propose scaling the network output and shifting the timestep, respectively. Inspired by these insights, we introduce two learnable parameters, $o_{n}$ and $\delta_{n}^{k}$, to perturb the scale of network output’s and the timestep. Our EPD-Solver follows the update rule:
@@
To determine ${\Theta}_{1:N}$, we employ a distillation-based optimization process. Specifically, given a student time schedule with $N$ steps $\mathcal{T}_{\mathsf{stu}}=\{t_{0}=t_{\min},...,t_{N}=t_{\max}\}$, we insert $M$ intermediate steps between $t_{n}$ and $t_{n+1}$, i.e., $\mathcal{T}_{\mathsf{tea}}=\{t_{0},...,t_{n},t_{n}^{1},...,t_{n}^{M},t_{n+1},..,t_{N}\}$, to yield a more accurate teacher trajectories. The training process starts with generating teacher trajectories by any ODE solver (e.g., DPM-Solver) and store the reference states as $\{{\mathbf{y}}_{t_{n}}\}_{n=0}^{N}$. Afterward, we sample student trajectory with the same initial noise ${\mathbf{y}}_{t_{N}}$, and optimize the parameters $\{\Theta_{n}\}_{n=1}^{N}$ to obtain the student trajectory $\{{\mathbf{x}}_{t_{n}}\}_{n=0}^{N}$ that aligns the teacher trajectory w.r.t some distance measurement $\text{dist}(\cdot,\cdot)$. For noisy states $\{{\mathbf{x}}_{t_{n}}\}_{n=1}^{N}$, we use the squared $\ell_{2}$ distance as $\text{dist}(\cdot,\cdot)$. For a generated sample ${\mathbf{x}}_{t_{0}}$, we compute the squared $\ell_{2}$ distance in the feature space of the last layer of an ImageNet-pretrained Inception network [szegedy2015going]. In particular, to improve the alignment between ${\mathbf{x}}_{t_{n}}$ and ${\mathbf{y}}_{t_{n}}$, since the value of ${\mathbf{x}}_{t_{n}}$ is dependent of the parameters $\Theta_{N}$ to $\Theta_{n}$, we aim to optimize them by minimizing
@@
In one training loop, we require $N$ backpropagation. The entire training algorithm is listed in algorithm 1 and the inference procedure is provided in algorithm 3. By default, we adopt the analytical first step (AFS) trick [dockhorn2022genie] in the first step to save one NFE by simply using ${\mathbf{x}}_{t_{N}}$ as direction.
@@
EPD-Plugin to existing solvers. EPD-Solver can be applied to existing solvers to further enhance diffusion sampling. The key idea is to replace their original gradient estimation with multiple parallel branches. As a representative case, we demonstrate this using the multi-step iPNDM sampler [liupseudo, zhangfast]. We refer to the modified solver as EPD-Plugin. Due to space limitations, a detailed description is deferred to section A-B.
@@ III-D Stage 2: Residual Dirichlet Policy Optimization
While existing learnable solvers predominantly rely on trajectory-preserving distillation to compress high-NFE ODE solvers [wang2025adaptivestochasticcoefficientsaccelerating, zhou2024fast, tong2024learning, chen2024trajectory], this paradigm often leads to noticeable degradation in the few-step regime. This limitation arises because, for extremely low-step solvers, learning an exact one-to-one mapping from noise to teacher trajectories is inherently challenging [tong2024learning], as fitting errors tend to accumulate along the shortened sampling path. Fortunately, for large-scale text-to-image diffusion models, human perception does not require strict numerical consistency with the teacher, but rather semantic and perceptual alignment. Motivated by these observations, we adopt a parameter-efficient reinforcement learning stage to move beyond exact trajectory matching and align the sampling behavior with human preferences.
@@
Policy optimization. In this paper, we adopt a lightweight PPO [schulman2017proximalpolicyoptimizationalgorithms] variant with reward leave-one-out (RLOO) advantages. Specifically, for each text prompt, we draw $G$ solvers from the policy and generate $G$ candidate images. Let $r_{g}$ denote the scalar reward of the $g$-th candidate, as computed by HPS v2.1 [hpsv2]. The RLOO reward baseline for $r_{g}$ is defined as the average reward of the other candidates from the same prompt:
@@ IV-A Setup
Models. We test out ODE solvers on diffusion-based image generation models, covering both pixel-space [karras2022elucidating] and latent-space models [rombach2022high], across image resolutions ranging from 32 to 1024. For pixel-space models, we evaluate the pretrained models on CIFAR $\times32$ [krizhevsky2009learning], FFHQ $\times64$ [karras2019style], ImageNet $\times64$ [russakovsky2015imagenet] from [karras2022elucidating]. For latent-space models, we evaluate pretrained models on LSUN Bedroom $\times256$ [yu2015lsun] from [rombach2022high], as well as large-scale text-to-image (T2I) models, including Stable Diffusion v1.5 [rombach2022high] at $\times512$ and SD3-Medium [sd3] at both $\times512$ and $\times1024$ resolutions.
Baseline solvers. We compare against representative ODE solvers across three categories: (1) Single-step solvers: DDIM [songdenoising], EDM [karras2022elucidating], DPM-Solver-2 [lu2022dpm], and AMED-Solver [zhou2024fast]; (2) Multi-step solvers: DPM-Solver++(3M)[lu2022dpm_plus], UniPC[zhao2024unipc], iPNDM [liupseudo, zhangfast], and AMED-Plugin [zhou2024fast]; (3) Parallelism-based solver: ParaDiGMS [shih2023parallel]. For a fair comparison, we follow the recommended time schedules from their original papers [karras2022elucidating, lu2022dpm_plus, zhao2024unipc]. Specifically, we use the logSNR schedule for DPM-Solver-2, DPM-Solver++(3M), and UniPC, the time-uniform schedule for AMED-Solver [zhou2024fast], while employing the polynomial time schedule with $\rho=7$ for the remaining baselines. Please refer to section A-C for implementation details of baseline solvers.
Evaluation. We evaluate our solvers with different sampling regimes corresponding to model scale. For validation experiments (non-T2I models), we operate under low-latency constraints ($\text{NFE}\in\{3,5,7,9\}$) with AFS [dockhorn2022genie] applied; for large-scale text-to-image (T2I) experiments, we adopt a 20-NFE setting. When $K=1$, our method shares the same computational cost as the baseline solvers. For $K>1$, although each step involves K1K-1 additional function evaluations, these evaluations are fully parallelizable and incur only minimal overhead in inference latency. We refer to the effective cost under parallel execution as Parallel NFE (Para. NFE).
We assess sample quality using the Fréchet Inception Distance (FID). For unconditional and class-conditional validation, FID is computed over 50k generated images. For T2I models, we evaluate FID by generating 10k images using prompts from the MS-COCO validation set [lin2014microsoft]. To further evaluate T2I alignment and human aesthetic preference, we evaluate on the DrawBench dataset [saharia2022photorealistic] with 1k generated images and report a series of metrics, including ImageReward [xu2023imagereward], HPS v2.1 [hpsv2], CLIP score [clip], and PickScore [pickscore].
@@
For T2I experiments, we further employ the residual Dirichlet policy framework (stage 2) that operates without a teacher model. We utilize both Stable Diffusion v1.5 [rombach2022high] and SD3-Medium [sd3] as backbones. The training is conducted on the Pick-a-Pic dataset [pickscore], employing HPSv2.1 [hpsv2] as the reward model to align with human preferences. For classifier-free guidance (CFG), we set the scale to 7.5 for Stable Diffusion v1.5 and 4.5 for SD3-Medium. Additional implementation details are provided in section A-A.
@@ IV-B Validation Experiments Results
Single-step DDIM [songdenoising] 93.36 49.66 27.93 18.43
Heun [karras2022elucidating] 306.2 97.67 37.28 15.76
DPM-Solver-2 [lu2022dpm] 155.7 57.30 10.20 4.98
AMED-Solver [zhou2024fast] 18.49 7.59 4.36 3.67
Multi-step DPM-Solver++(3M) [lu2022dpm_plus] 110.0 24.97 6.74 3.42
UniPC [zhao2024unipc] 109.6 23.98 5.83 3.21
iPNDM [liupseudo, zhangfast] 47.98 13.59 5.08 3.17
AMED-Plugin [zhou2024fast] 10.81 6.61 3.65 2.63
Parallel ParaDiGMS [shih2023parallel] 51.03 18.96 7.18 6.19
@@
Single-step DDIM [songdenoising] 78.21 43.93 28.86 21.01
Heun [karras2022elucidating] 356.5 116.7 54.51 28.86
DPM-Solver-2 [lu2022dpm] 266.0 87.10 22.59 9.26
AMED-Solver [zhou2024fast] 47.31 14.80 8.82 6.31
Multi-step DPM-Solver++(3M) [lu2022dpm_plus] 86.45 22.51 8.44 4.77
UniPC [zhao2024unipc] 86.43 21.40 7.44 4.47
iPNDM [liupseudo, zhangfast] 45.98 17.17 7.79 4.58
AMED-Plugin [zhou2024fast] 26.87 12.49 6.64 4.24
Parallel ParaDiGMS [shih2023parallel] 43.64 20.92 16.39 8.81
@@
Single-step DDIM [songdenoising] 82.96 43.81 27.46 19.27
Heun [karras2022elucidating] 249.4 89.63 37.65 16.76
DPM-Solver-2 [lu2022dpm] 140.2 42.41 12.03 6.64
AMED-Solver [zhou2024fast] 38.10 10.74 6.66 5.44
Multi-step DPM-Solver++(3M) [lu2022dpm_plus] 91.52 25.49 10.14 6.48
UniPC [zhao2024unipc] 91.38 24.36 9.57 6.34
iPNDM [liupseudo, zhangfast] 58.53 18.99 9.17 5.91
AMED-Plugin [zhou2024fast] 28.06 13.83 7.81 5.60
Parallel ParaDiGMS [shih2023parallel] 41.11 17.27 13.67 6.38
@@
Single-step DDIM [songdenoising] 86.13 34.34 19.50 13.26
Heun [karras2022elucidating] 291.5 175.7 78.67 35.67
DPM-Solver-2 [lu2022dpm] 210.6 80.60 23.25 9.61
AMED-Solver [zhou2024fast] 58.21 13.20 7.10 5.65
Multi-step DPM-Solver++(3M) [lu2022dpm_plus] 111.9 23.15 8.87 6.45
UniPC [zhao2024unipc] 112.3 23.34 8.73 6.61
iPNDM [liupseudo, zhangfast] 80.99 26.65 13.80 8.38
AMED-Plugin [zhou2024fast] 101.5 25.68 8.63 7.82
Parallel ParaDiGMS [shih2023parallel] 100.3 31.68 15.85 8.56
@@ IV-C Text-to-Image Experiment Results
TABLE II: Quantitative comparison of solvers on Stable Diffusion v1.5 (512 $\times$ 512) [rombach2022high]. The best results are in bold, and the second best are underlined. Qualitative results are in Figure 5. See Table XVII (a) for the value of the learned parameters.
@@
DDIM (Default) 50 Uniform 0.2454 21.1705 3.0980 0.2701 5.2630
iPNDM (3M) 50 Uniform 0.2474 21.1860 3.1234 0.2700 5.2599
Heun 25 Polynomial 0.2471 21.1498 3.1167 0.2697 5.2385
Uniform 0.2473 21.1733 3.1229 0.2701 5.2522
DPM-Solver-2 25 Uniform 0.2473 21.1632 3.1263 0.2699 5.2429
LogSNR 0.2470 21.1632 3.1275 0.2702 5.2411
@@
DDIM 20 Uniform 0.2416 21.1275 3.0589 0.2705 5.2552
iPNDM (3M) 20 Uniform 0.2454 21.1341 3.0830 0.2691 5.2546
Heun 10 Polynomial 0.2387 20.8721 2.9216 0.2680 5.1193
Uniform 0.2443 21.0318 3.0756 0.2692 5.1723
DPM-Solver-2 10 LogSNR 0.2364 20.7718 2.9378 0.2660 5.1278
Uniform 0.2435 20.9930 3.0796 0.2689 5.1708
EPD-Solver 10 Uniform 0.2482 21.1302 3.1121 0.2695 5.2388
TABLE III: Quantitative comparison of solvers on SD3-Medium (512 $\times$ 512) [sd3]. The best results are in bold, and the second best are underlined. Qualitative results are in Figure 18. See Table XVII (b) for the value of the learned parameters.
@@
DDIM (Default) 28 Shift Time 0.2734 22.0357 3.8877 0.2820 5.2562
Heun 14 Shift Time 0.2685 21.8897 3.8622 0.2819 5.1382
DPM-Solver-2 14 Shift Time 0.2656 21.8513 3.8367 0.2810 5.1047
iPNDM (3M) 28 Shift Time 0.2700 21.8742 3.8624 0.2800 5.1802
@@
DDIM 20 Shift Time 0.2707 21.9684 3.8585 0.2807 5.2620
Heun 10 Shift Time 0.2644 21.8026 3.8095 0.2814 5.0824
DPM-Solver-2 10 Shift Time 0.2631 21.7358 3.8081 0.2798 5.0572
iPNDM (3M) 20 Shift Time 0.2690 21.7928 3.8467 0.2799 5.1692
EPD-Solver 10 Shift Time 0.2742 21.9514 3.8856 0.2813 5.2743
TABLE IV: Quantitative comparison of solvers on SD3-Medium (1024 $\times$ 1024) [sd3]. The best results are in bold, and the second best are underlined. Qualitative results are in Figure 19. See Table XVII (c) for the value of the learned parameters.
@@
DDIM (Default) 28 Shift Time 0.2820 22.4839 3.9796 0.2854 5.3689
Heun 14 Shift Time 0.2790 22.3832 3.9622 0.2853 5.2688
DPM-Solver-2 14 Shift Time 0.2817 22.4119 4.0027 0.2866 5.2590
iPNDM (3M) 28 Shift Time 0.2818 22.4841 4.0057 0.2850 5.3565
@@
DDIM 20 Shift Time 0.2769 22.3774 3.9240 0.2850 5.3623
Heun 10 Shift Time 0.2707 22.2173 3.8767 0.2871 5.1892
DPM-Solver-2 10 Shift Time 0.2759 22.2812 3.9323 0.2874 5.2458
iPNDM (3M) 20 Shift Time 0.2805 22.3740 3.9633 0.2840 5.3460
EPD-Solver 10 Shift Time 0.2823 22.3942 3.9765 0.2852 5.3995
@@
Table II compares the quantitative results on Stable Diffusion v1.5 [rombach2022high]. Under the low-latency setting (NFE = 20), EPD-Solver demonstrates superior human alignment, achieving an HPSv2.1 score of 0.2482 [hpsv2]. This result not only surpasses all baselines at the same computational budget but also outperforms the best-performing method at NFE=50 (iPNDM, 0.2474) [zhangfast]. Furthermore, our RL fine-tuning significantly boosts the distilled base model, improving the ImageReward from 3.0405 to 3.1121 [xu2023imagereward]. Overall, EPD-Solver achieves generation quality competitive with 50-step solvers while requiring only 40% of the inference steps, effectively bridging the gap between efficiency and generation fidelity.
We further validate the scalability of our method on the recent SD3-Medium [sd3] across different resolutions. As shown in Table III, at $512\times 512$ resolution with 20 NFE, our RL-tuned solver achieves an HPSv2.1 score of 0.2742, outperforming the offical setting (DDIM at 28 NFE, 0.2734) in terms of human preference alignment. While the ImageReward score (3.8856) is slightly lower than the 28-step baseline (3.8877), it remains highly competitive given the reduced computational cost. Similarly, at $1024\times 1024$ resolution (Table IV), EPD-Solver maintains its advantage with an HPSv2.1 score of 0.2823, surpassing the 28-step DDIM baseline (0.2820). To examine the training dynamics, Figure 7 plots the HPS v2.1 curves, demonstrating rapid convergence and consistent superiority over baselines across all resolutions. Complementing this, Figure 6 visualizes the qualitative evolution of generated samples. These findings indicate that our Residual Dirichlet Policy Optimization generalizes effectively to large-scale text-to-image generation.

### v2

@@ III-B1 Discussion with multi-step solvers
While multi-step solvers [35, 76, 32, 71] also combine multiple gradients to approximate the integral, they fundamentally differ in where these gradients are evaluated. Specifically, multi-step methods rely on Taylor expansion or polynomial extrapolation to linearly combine historical gradients evaluated at previous time steps, i.e., outside the current integration interval. In contrast, Theorem 1 implies that the exact integral over an interval $[a,b]$ admits a representation as a convex combination of gradients evaluated at points strictly within $(a,b)$. Motivated by this theoretical result, our method explicitly constructs a simplex-weighted combination of multiple gradients evaluated within the current time interval, leading to a more faithful approximation of the integral.
@@ III-B2 Discussion with AMED-Solver
AMED-Solver [80] estimates the update direction using a single intermediate timestep. However, according to Theorem 1, the integral of a vector-valued function cannot, in general, be exactly represented by the derivative at a single timestep; instead, it admits a convex combination of derivatives evaluated at multiple timesteps. A single timestep may only suffice when the underlying trajectory is effectively one-dimensional. In the sequel, we empirically analyze the geometric properties of diffusion sampling trajectories and find that they are nearly confined to a two-dimensional manifold, violating this condition.
In Figure 3, we analyze diffusion sampling trajectories using the models from [16]. Despite the high dimensionality of the ambient space, the cumulative explained variance ratio shows that over 97% of the residual variance is captured by the first two principal components, indicating that the non-linear trajectory is effectively confined to a two-dimensional manifold. As a result, single-step or single-intermediate-point methods such as AMED-Solver are generally insufficient to characterize the local curvature within this plane using a single direction. This motivates the use of multiple intermediate gradients to span the underlying subspace and achieve a more accurate integral approximation.
@@ III-C Stage 1: Distillation-Based Parameter Optimization
[42, 25] identify exposure bias—i.e., the mismatch between training and sampling inputs—as a key factor contributing to error accumulation and sampling drift. To mitigate this, they propose scaling the network output and shifting the timestep, respectively. Inspired by these insights, we introduce two learnable parameters, $o_{n}$ and $\delta_{n}^{k}$, to perturb the scale of network output’s and the timestep. Our EPD-Solver follows the update rule:
@@
To determine ${\Theta}_{1:N}$, we employ a distillation-based optimization process. Specifically, given a student time schedule with $N$ steps $\mathcal{T}_{\mathsf{stu}}=\{t_{0}=t_{\min},...,t_{N}=t_{\max}\}$, we insert $M$ intermediate steps between $t_{n}$ and $t_{n+1}$, i.e., $\mathcal{T}_{\mathsf{tea}}=\{t_{0},...,t_{n},t_{n}^{1},...,t_{n}^{M},t_{n+1},..,t_{N}\}$, to yield a more accurate teacher trajectories. The training process starts with generating teacher trajectories by any ODE solver (e.g., DPM-Solver) and store the reference states as $\{{\mathbf{y}}_{t_{n}}\}_{n=0}^{N}$. Afterward, we sample student trajectory with the same initial noise ${\mathbf{y}}_{t_{N}}$, and optimize the parameters $\{\Theta_{n}\}_{n=1}^{N}$ to obtain the student trajectory $\{{\mathbf{x}}_{t_{n}}\}_{n=0}^{N}$ that aligns the teacher trajectory w.r.t some distance measurement $\text{dist}(\cdot,\cdot)$. For noisy states $\{{\mathbf{x}}_{t_{n}}\}_{n=1}^{N}$, we use the squared $\ell_{2}$ distance as $\text{dist}(\cdot,\cdot)$. For a generated sample ${\mathbf{x}}_{t_{0}}$, we compute the squared $\ell_{2}$ distance in the feature space of the last layer of an ImageNet-pretrained Inception network [58]. In particular, to improve the alignment between ${\mathbf{x}}_{t_{n}}$ and ${\mathbf{y}}_{t_{n}}$, since the value of ${\mathbf{x}}_{t_{n}}$ is dependent of the parameters $\Theta_{N}$ to $\Theta_{n}$, we aim to optimize them by minimizing
@@
In one training loop, we require $N$ backpropagation. The entire training algorithm is listed in algorithm 1 and the inference procedure is provided in algorithm 3. By default, we adopt the analytical first step (AFS) trick [7] in the first step to save one NFE by simply using ${\mathbf{x}}_{t_{N}}$ as direction.
@@
EPD-Plugin to existing solvers. EPD-Solver can be applied to existing solvers to further enhance diffusion sampling. The key idea is to replace their original gradient estimation with multiple parallel branches. As a representative case, we demonstrate this using the multi-step iPNDM sampler [32, 71]. We refer to the modified solver as EPD-Plugin. Due to space limitations, a detailed description is deferred to section A-B.
@@ III-D Stage 2: Residual Dirichlet Policy Optimization
While existing learnable solvers predominantly rely on trajectory-preserving distillation to compress high-NFE ODE solvers [61, 80, 59, 4], this paradigm often leads to noticeable degradation in the few-step regime. This limitation arises because, for extremely low-step solvers, learning an exact one-to-one mapping from noise to teacher trajectories is inherently challenging [59], as fitting errors tend to accumulate along the shortened sampling path. Fortunately, for large-scale text-to-image diffusion models, human perception does not require strict numerical consistency with the teacher, but rather semantic and perceptual alignment. Motivated by these observations, we adopt a parameter-efficient reinforcement learning stage to move beyond exact trajectory matching and align the sampling behavior with human preferences.
@@
Policy optimization. In this paper, we adopt a lightweight PPO [52] variant with reward leave-one-out (RLOO) advantages. Specifically, for each text prompt, we draw $G$ solvers from the policy and generate $G$ candidate images. Let $r_{g}$ denote the scalar reward of the $g$-th candidate, as computed by HPS v2.1 [64]. The RLOO reward baseline for $r_{g}$ is defined as the average reward of the other candidates from the same prompt:
@@ IV-A Setup
Models. We test out ODE solvers on diffusion-based image generation models, covering both pixel-space [15] and latent-space models [46], across image resolutions ranging from 32 to 1024. For pixel-space models, we evaluate the pretrained models on CIFAR $\times32$ [22], FFHQ $\times64$ [17], ImageNet $\times64$ [47] from [15]. For latent-space models, we evaluate pretrained models on LSUN Bedroom $\times256$ [70] from [46], as well as large-scale text-to-image (T2I) models, including Stable Diffusion v1.5 [46] at $\times512$ and SD3-Medium [9] at both $\times512$ and $\times1024$ resolutions.
Baseline solvers. We compare against representative ODE solvers across three categories: (1) Single-step solvers: DDIM [55], EDM [15], DPM-Solver-2 [34], and AMED-Solver [80]; (2) Multi-step solvers: DPM-Solver++(3M)[35], UniPC[76], iPNDM [32, 71], and AMED-Plugin [80]; (3) Parallelism-based solver: ParaDiGMS [53]. For a fair comparison, we follow the recommended time schedules from their original papers [15, 35, 76]. Specifically, we use the logSNR schedule for DPM-Solver-2, DPM-Solver++(3M), and UniPC, the time-uniform schedule for AMED-Solver [80], while employing the polynomial time schedule with $\rho=7$ for the remaining baselines. Please refer to section A-C for implementation details of baseline solvers.
Evaluation. We evaluate our solvers with different sampling regimes corresponding to model scale. For validation experiments (non-T2I models), we operate under low-latency constraints ($\text{NFE}\in\{3,5,7,9\}$) with AFS [7] applied; for large-scale text-to-image (T2I) experiments, we adopt a 20-NFE setting. When $K=1$, our method shares the same computational cost as the baseline solvers. For $K>1$, although each step involves K1K-1 additional function evaluations, these evaluations are fully parallelizable and incur only minimal overhead in inference latency. We refer to the effective cost under parallel execution as Parallel NFE (Para. NFE).
We assess sample quality using the Fréchet Inception Distance (FID). For unconditional and class-conditional validation, FID is computed over 50k generated images. For T2I models, we evaluate FID by generating 10k images using prompts from the MS-COCO validation set [28]. To further evaluate T2I alignment and human aesthetic preference, we evaluate on the DrawBench dataset [49] with 1k generated images and report a series of metrics, including ImageReward [65], HPS v2.1 [64], CLIP score [45], and PickScore [21].
@@
For T2I experiments, we further employ the residual Dirichlet policy framework (stage 2) that operates without a teacher model. We utilize both Stable Diffusion v1.5 [46] and SD3-Medium [9] as backbones. The training is conducted on the Pick-a-Pic dataset [21], employing HPSv2.1 [64] as the reward model to align with human preferences. For classifier-free guidance (CFG), we set the scale to 7.5 for Stable Diffusion v1.5 and 4.5 for SD3-Medium. Additional implementation details are provided in section A-A.
@@ IV-B Validation Experiments Results
Single-step DDIM [55] 93.36 49.66 27.93 18.43
Heun [15] 306.2 97.67 37.28 15.76
DPM-Solver-2 [34] 155.7 57.30 10.20 4.98
AMED-Solver [80] 18.49 7.59 4.36 3.67
Multi-step DPM-Solver++(3M) [35] 110.0 24.97 6.74 3.42
UniPC [76] 109.6 23.98 5.83 3.21
iPNDM [32, 71] 47.98 13.59 5.08 3.17
AMED-Plugin [80] 10.81 6.61 3.65 2.63
Parallel ParaDiGMS [53] 51.03 18.96 7.18 6.19
@@
Single-step DDIM [55] 78.21 43.93 28.86 21.01
Heun [15] 356.5 116.7 54.51 28.86
DPM-Solver-2 [34] 266.0 87.10 22.59 9.26
AMED-Solver [80] 47.31 14.80 8.82 6.31
Multi-step DPM-Solver++(3M) [35] 86.45 22.51 8.44 4.77
UniPC [76] 86.43 21.40 7.44 4.47
iPNDM [32, 71] 45.98 17.17 7.79 4.58
AMED-Plugin [80] 26.87 12.49 6.64 4.24
Parallel ParaDiGMS [53] 43.64 20.92 16.39 8.81
@@
Single-step DDIM [55] 82.96 43.81 27.46 19.27
Heun [15] 249.4 89.63 37.65 16.76
DPM-Solver-2 [34] 140.2 42.41 12.03 6.64
AMED-Solver [80] 38.10 10.74 6.66 5.44
Multi-step DPM-Solver++(3M) [35] 91.52 25.49 10.14 6.48
UniPC [76] 91.38 24.36 9.57 6.34
iPNDM [32, 71] 58.53 18.99 9.17 5.91
AMED-Plugin [80] 28.06 13.83 7.81 5.60
Parallel ParaDiGMS [53] 41.11 17.27 13.67 6.38
@@
Single-step DDIM [55] 86.13 34.34 19.50 13.26
Heun [15] 291.5 175.7 78.67 35.67
DPM-Solver-2 [34] 210.6 80.60 23.25 9.61
AMED-Solver [80] 58.21 13.20 7.10 5.65
Multi-step DPM-Solver++(3M) [35] 111.9 23.15 8.87 6.45
UniPC [76] 112.3 23.34 8.73 6.61
iPNDM [32, 71] 80.99 26.65 13.80 8.38
AMED-Plugin [80] 101.5 25.68 8.63 7.82
Parallel ParaDiGMS [53] 100.3 31.68 15.85 8.56
@@ IV-C Text-to-Image Experiment Results
TABLE II: Quantitative comparison of solvers on Stable Diffusion v1.5 (512 $\times$ 512) [46]. The best results are in bold, and the second best are underlined. Qualitative results are in Figure 5. See Table XVII (a) for the value of the learned parameters.
@@
DDIM (Default) 50 Uniform 0.2454 21.1705 -0.0020 0.2701 5.2630
iPNDM (3M) 50 Uniform 0.2474 21.1860 0.0234 0.2700 5.2599
Heun 25 Polynomial 0.2471 21.1498 0.0167 0.2697 5.2385
Uniform 0.2473 21.1733 0.0229 0.2701 5.2522
DPM-Solver-2 25 Uniform 0.2473 21.1632 0.0263 0.2699 5.2429
LogSNR 0.2470 21.1632 0.0275 0.2702 5.2411
@@
DDIM 20 Uniform 0.2416 21.1275 -0.0411 0.2705 5.2552
iPNDM (3M) 20 Uniform 0.2454 21.1341 -0.0170 0.2691 5.2546
Heun 10 Polynomial 0.2387 20.8721 -0.1784 0.2680 5.1193
Uniform 0.2443 21.0318 -0.0244 0.2692 5.1723
DPM-Solver-2 10 LogSNR 0.2364 20.7718 -0.1622 0.2660 5.1278
Uniform 0.2435 20.9930 -0.0204 0.2689 5.1708
EPD-Solver 10 Uniform 0.2482 21.1302 0.0121 0.2695 5.2388
TABLE III: Quantitative comparison of solvers on SD3-Medium (512 $\times$ 512) [9]. The best results are in bold, and the second best are underlined. Qualitative results are in Figure 18. See Table XVII (b) for the value of the learned parameters.
@@
DDIM (Default) 28 Shift Time 0.2734 22.0357 0.7877 0.2820 5.2562
Heun 14 Shift Time 0.2685 21.8897 0.7622 0.2819 5.1382
DPM-Solver-2 14 Shift Time 0.2656 21.8513 0.7367 0.2810 5.1047
iPNDM (3M) 28 Shift Time 0.2700 21.8742 0.7624 0.2800 5.1802
@@
DDIM 20 Shift Time 0.2707 21.9684 0.7585 0.2807 5.2620
Heun 10 Shift Time 0.2644 21.8026 0.7095 0.2814 5.0824
DPM-Solver-2 10 Shift Time 0.2631 21.7358 0.7081 0.2798 5.0572
iPNDM (3M) 20 Shift Time 0.2690 21.7928 0.7467 0.2799 5.1692
EPD-Solver 10 Shift Time 0.2742 21.9514 0.7856 0.2813 5.2743
TABLE IV: Quantitative comparison of solvers on SD3-Medium (1024 $\times$ 1024) [9]. The best results are in bold, and the second best are underlined. Qualitative results are in Figure 19. See Table XVII (c) for the value of the learned parameters.
@@
DDIM (Default) 28 Shift Time 0.2820 22.4839 0.8796 0.2854 5.3689
Heun 14 Shift Time 0.2790 22.3832 0.8622 0.2853 5.2688
DPM-Solver-2 14 Shift Time 0.2817 22.4119 0.9027 0.2866 5.2590
iPNDM (3M) 28 Shift Time 0.2818 22.4841 0.9057 0.2850 5.3565
@@
DDIM 20 Shift Time 0.2769 22.3774 0.8240 0.2850 5.3623
Heun 10 Shift Time 0.2707 22.2173 0.7767 0.2871 5.1892
DPM-Solver-2 10 Shift Time 0.2759 22.2812 0.8323 0.2874 5.2458
iPNDM (3M) 20 Shift Time 0.2805 22.3740 0.8633 0.2840 5.3460
EPD-Solver 10 Shift Time 0.2823 22.3942 0.8765 0.2852 5.3995
@@
Table II compares the quantitative results on Stable Diffusion v1.5 [46]. Under the low-latency setting (NFE = 20), EPD-Solver demonstrates superior human alignment, achieving an HPSv2.1 score of 0.2482 [64]. This result not only surpasses all baselines at the same computational budget but also outperforms the best-performing method at NFE=50 (iPNDM, 0.2474) [71]. Furthermore, our RL fine-tuning significantly boosts the distilled base model, improving the ImageReward from 3.0405 to 3.1121 [65]. Overall, EPD-Solver achieves generation quality competitive with 50-step solvers while requiring only 40% of the inference steps, effectively bridging the gap between efficiency and generation fidelity.
We further validate the scalability of our method on the recent SD3-Medium [9] across different resolutions. As shown in Table III, at $512\times 512$ resolution with 20 NFE, our RL-tuned solver achieves an HPSv2.1 score of 0.2742, outperforming the offical setting (DDIM at 28 NFE, 0.2734) in terms of human preference alignment. While the ImageReward score (3.8856) is slightly lower than the 28-step baseline (3.8877), it remains highly competitive given the reduced computational cost. Similarly, at $1024\times 1024$ resolution (Table IV), EPD-Solver maintains its advantage with an HPSv2.1 score of 0.2823, surpassing the 28-step DDIM baseline (0.2820). To examine the training dynamics, Figure 7 plots the HPS v2.1 curves, demonstrating rapid convergence and consistent superiority over baselines across all resolutions. Complementing this, Figure 6 visualizes the qualitative evolution of generated samples. These findings indicate that our Residual Dirichlet Policy Optimization generalizes effectively to large-scale text-to-image generation.